# Profiling stellar environments of gravitational wave sources

### v1

@@ Abstract
Introduction.— A number of efforts trying to probe the provenance of compact binary coalescences (CBCs) detected with the LIGO-Virgo-KAGRA (LVK) gravitational wave (GW) detector network Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021); Abbott et al. (2023) are currently underway (see, e.g., Mapelli (2021) for a review). Indeed, determining the environment in which the CBC formed, evolved, and merged is highly non-trivial for two primary reasons. First, the GW waveform pertaining to the CBC does not always contain signatures of the CBC’s formation channel. While statistical arguments based on estimates of the CBC’s intrinsic source parameters could provide hints of its host environment, these cannot in general conclusively rule out all other possible environments in favor of the statistically preferred one Zevin et al. (2021); Pierra et al. (2024). Second, the lack of observed electromagnetic counterparts for all but one of the detected CBC events Abbott et al. (2023), in tandem with poor localisation sky-areas spanning $\mathcal{O}(10)$ sq. deg. or larger, make it difficult to identify the host via direct observations Chen and Holz (2016). The inclusion of additional detectors in current and forthcoming observing runs, such as KAGRA and LIGO-India Saleem et al. (2022), as well as planned and proposed next-generation ground and space-based missions, will have markedly improved localisation sky-areas, even going below $\mathcal{O}(1)$ sq. deg. for space-based missions Zhang et al. (2021). However, confident identification of the CBC’s host will require arc- or sub-arc-second localisations, which will in general not be achievable even by future detectors, especially at large redshifts.
@@
Kinematic parameters (such as acceleration) associated with the motion of the center of mass (CoM) of the CBC, could serve as messengers of the CBC’s environment. Past work has shown that a finite line-of-sight acceleration of the CBC’s CoM will modulate the GW waveform with respect to one produced by a non-accelerated CBC  Vijaykumar et al. (2023); Bonvin et al. (2017); Yunes et al. (2011). At leading order, these modulations occur at $-4$ PN (Post Newtonian order). Higher time derivatives of this acceleration (e.g. jerk, snap, etc) cause modulations that start at even lower PN orders (e.g. $-8$ PN, $-16$ PN, etc) Yunes et al. (2011). Thus, detectors with sensitivities at low frequencies, such as space-based ones, are best suited to extract these kinematic parameters, although sufficiently large accelerations might be observable even with ground-based detectors Vijaykumar et al. (2023).
@@
Most proposed formation channels of stellar-mass CBCs suggest that these either underwent isolated evolution in the galactic field, or were assembled dynamically in dense stellar environments such as globular clusters (GCs), nuclear star clusters (NSCs), and active galactic nuclei (AGNs). These environments have characteristic properties, in particular, mass and potential profiles, which govern the kinematics of the CoMs of the CBCs they host, thereby imprinting themselves on the emanated GWs. In this Letter, we demonstrate how these properties can be directly extracted from the CBC’s GWs, thus effectively identifying the CBC’s environment on a single-event basis.
@@
Method.— The total mass of a binary $M=m_{1}+m_{2}$ in the detector frame as inferred from the GW signal is related to the source frame total mass $M_{\mathrm{src}}$ as:
@@
where $z_{\mathrm{cos}}$ is the cosmological redshift, and $z_{\mathrm{dop}}$ is the Doppler shift caused by any line of sight velocity (LOSV) of the CoM of the binary [^1] From hereon, $M$ will denote the detector frame total mass.. A constant LOSV is fully degenerate with the source frame mass. However, a time-varying LOSV results in a time-varying Doppler shift and induces a time-varying detector frame mass, allowing time derivatives of the LOSV to be measured. Modulations to the GW waveform appear at different Post-Newtonian (PN) orders depending on the order of the time derivative of LOSV causing the modulation.
$M_{z}=M\left(1+\sum_{n=1}^{\infty}\Gamma_{n}(t_{o}-t_{c})^{n}\right)$ (2)
where $\Gamma_{n}\equiv\left.\frac{1}{c}\frac{1}{n!}\frac{d^{n}v_{l}}{dt_{o}^{n}}\right|_{t_{o}=t_{c}}$, $t_{o}$ is the observation time, $t_{c}$ is the time at coalescence, and $c$ is the speed of light. Note that Eq. (2) is valid under the assumption that $|\Gamma_{n}(t_{o}-t_{c})^{n}|$ are $\ll$ 1222As a consistency check of this condition, we ensure that $|\Gamma_{n}(t_{o}-t_{c})^{n}|\leq 0.05$., where special relativistic effects may be neglected.
The corresponding modulated GW waveform, in the stationary phase approximation, may be written as Buonanno et al. (2009):
$\tilde{h}(f)=\mathcal{A}f^{-7/6}e^{i\left(\Psi(f)+\Delta\Psi_{-4n}(f)\right)}$ (3)
where $\mathcal{A}\propto\frac{\mathcal{M}^{5/6}}{D_{L}}$ is the amplitude, $\mathcal{M}$ is the chirp mass, $D_{L}$ is the luminosity distance, $\Psi(f)$ is the unperturbed phase given by Eq. 3.18 of Buonanno et al. (2009), and $\Delta\Psi_{-4n}(f)$ is a $-4n$ PN phase correction due to $n^{th}$ derivative of $v_{l}$. At leading order, this is given by Tiwari et al. (2024a):
@@
where $v=(\pi Mf)^{1/3}$, $\eta\equiv m_{1}m_{2}/(m_{1}+m_{2})^{2}$ is the symmetric mass ratio, and $f$ is the observed GW frequency.
Positive powers of a kinematic parameter can contribute to PN orders lower than the corresponding leading order. For example, the $-8$PN phase correction $\Delta\Psi_{-8}(f)$ will have contributions from the line of sight jerk (LOSJ) $\Gamma_{2}$ of the CoM, as well as the LOSA term proportional to $\Gamma_{1}^{2}$. We account for these corrections as well, down to $-24$ PN, below which we do not expect PN terms to be well constrained even with detectors sensitive down to mili-hertz frequencies [^3] See Figure 6 and the corresponding section in the Supplemental Material Tiwari et al. (2024a)..
@@
Figure 2: Circular outer-orbit around an SMBH: The precision of the extraction of parameters over a grid of SMBH masses $M_{\rm SMBH}$ and radii of outer orbit $R$ for a binary system of component masses (from left to right): $5-1.4\,M_{\odot}$ at $\mathrm{100Mpc}$ in A+, $30-30\,M_{\odot}$ at $100\mathrm{Mpc}$ in ET, and $100-100\,M_{\odot}$ at $1\mathrm{Gpc}$ in DECIGO and LISA, with $\cos\vartheta=0.9$. Top Panels: Relative error in the extraction of mass of SMBHs. The ‘solid’ lines are the lines of constant $\frac{\Delta M_{\rm SMBH}}{M_{\rm SMBH}}$. Bottom Panel: Relative error in the extraction of radii of the outer orbit. The ‘dashed’ lines are the lines of constant $\frac{\Delta R}{R}$. The pink regions correspond to the parameter space where either the relative errors are greater than 1 or we do not sample them due to the non-measurability of the kinematic parameters. The blue lines demarcate the regions (lower left) where we do not perform the Fisher matrix analysis, for, in these regions, the assumption $|\Gamma_{n}(t_{o}-t_{c})^{n}|\ll 1$ is no longer valid. Note that the arrows point towards the region of increasing precision in the inference of parameters.
Figure 3: Eccentric outer-orbit around an SMBH: The precision of the extraction of parameters over a grid of SMBH masses $M_{\rm SMBH}$ and semi-major axes $R$ of the outer orbit for a binary system of component masses (from left to right) considered in Figure 2 with $e=0.5$ and $x\approx 1.25$. $x$ was drawn randomly from a uniform distribution between $\frac{1}{1-e}$ and $\frac{1}{1+e}$. Top Panel: Relative error in the extraction of mass of SMBHs. The ‘solid’ lines are the lines of constant $\frac{\Delta M_{\rm SMBH}}{M_{\rm SMBH}}$. Bottom Panel: Relative error in the extraction of the location of binary in the outer orbit. The ‘dotted’ lines are the lines of constant $\frac{\Delta x}{x}$. The pink regions, the blue lines, and the arrows have the same meaning as in Figure 2.
We consider three environments that could impart time-varying LOSVs to a CBC: A) the vicinity of a supermassive black hole (SMBH) as a proxy for an AGN environment, B) a globular cluster (GC), and C) a nuclear star cluster (NSC). To analyse the kinematics of the CoM of the CBC in these environments, we set up a Cartesian coordinate system in Figure 1 placed at the center of the gravitating source. Let $\iota_{\mathrm{out}}$ be the angle between our LOS and the angular momentum vector of the outer orbit, and let the projection vector of our LOS onto the outer orbital plane be along $\bm{e}_{y}$. Then:
@@
where $\cdot\equiv\frac{d}{dt}$ and ${}^{\prime}\equiv\frac{d}{dr}$. Using this relation iteratively, we calculate different kinematic parameters of the CoM of the binary and then project them onto our LOS ($\bm{n}$). This LOS component of the kinematic parameter will be a function of environment and orbit parameters, such as the steepness of the gravitational potential profile, and the location of the CBC with respect to the center of the potential, at coalescence [^4] For instance, a circular orbit of radius $R$ in a Keplerian potential around a SMBH of mass $M_{\rm SMBH}$ would have $\Gamma_{n}\sim\Omega^{n+1}R$, where $\Omega^{2}\sim M_{\rm SMBH}/R^{3}$.. We provide the derivation as well as the mathematical expressions of the LOS components in the Supplementary Material Tiwari et al. (2024a).
@@
To assess our ability to constrain these parameters in various observing scenarios, we adopt a Fisher Matrix approach Cutler and Flanagan (1994). The PN corrections to the GW phase due to the kinematic parameters are rewritten in terms of the environment and outer-orbit parameters. A Fisher Matrix on these is constructed and then inverted to yield a Covariance Matrix, which provides the Gaussian approximation to the GW likelihood. The environment parameters are sampled from this likelihood using emcee Foreman-Mackey et al. (2013), and the relative r.m.s errors on these parameters are estimated. We delineate these steps in detail in the Supplementary Material Tiwari et al. (2024a).

### v2

@@ I Introduction
A number of efforts trying to probe the provenance of compact binary coalescences (CBCs) detected with the LIGO-Virgo-KAGRA (LVK) gravitational wave (GW) detector network Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021); Abbott et al. (2023) are currently underway (see, e.g., Mapelli (2021) for a review). Indeed, determining the environment in which the CBC formed, evolved, and merged is highly non-trivial for two primary reasons. First, the GW waveform pertaining to the CBC does not always contain signatures of the CBC’s formation channel. While statistical arguments based on estimates of the CBC’s intrinsic source parameters could provide hints of its host environment, these cannot in general conclusively rule out all other possible environments in favor of the statistically preferred one Zevin et al. (2021); Pierra et al. (2024). Second, the lack of observed EM counterparts for all but one of the detected CBC events Abbott et al. (2017a, b, 2023), in tandem with poor localisation sky-areas spanning $\mathcal{O}(10)$ sq. deg. or larger, make it difficult to identify the host via direct observations Chen and Holz (2016). The inclusion of additional detectors in current and forthcoming observing runs, such as KAGRA and LIGO-India Saleem et al. (2022), as well as planned and proposed next-generation ground and space-based missions, will have markedly improved localisation sky-areas, even going below $\mathcal{O}(1)$ sq. deg. for space-based missions Zhang et al. (2021). However, confident identification of the CBC’s host will require arc- or sub-arc-second localisations, which will in general not be achievable even by future detectors, especially at large redshifts.
@@
Kinematic parameters (such as acceleration) associated with the motion of the center of mass (CoM) of the CBC, could serve as messengers of the CBC’s environment. Past work has shown that a finite line-of-sight acceleration of the CBC’s CoM will modulate the GW waveform with respect to one produced by a non-accelerated CBC Vijaykumar et al. (2023); Bonvin et al. (2017); Yunes et al. (2011). At leading order, these modulations occur at 4-4 post Newtonian (PN) order. Higher time derivatives of this acceleration (e.g. jerk, snap, etc) cause modulations that start at even lower PN orders (e.g. 8-8 PN, 12-12 PN, etc) Yunes et al. (2011). Thus, detectors with sensitivities at low frequencies, such as space-based ones, are best suited to extract these kinematic parameters, although sufficiently large accelerations might be observable even with ground-based detectors Vijaykumar et al. (2023).
@@
Most proposed formation channels of stellar-mass CBCs suggest that these either underwent isolated evolution in the galactic field, or were assembled dynamically in dense stellar environments such as globular clusters (GCs), nuclear star clusters (NSCs), and active galactic nuclei (AGNs). These environments have characteristic properties, in particular, mass and potential profiles, which govern the kinematics of the CoMs of the CBCs they host, thereby imprinting themselves on the emanated GWs. In this paper, we demonstrate how these properties can be directly extracted from the CBC’s GWs, thus effectively identifying the CBC’s environment on a single-event basis.
@@
II Method
II.1 Phase Correction
The total mass of a binary $M=m_{1}+m_{2}$ in the detector frame as inferred from the GW signal is related to the source frame total mass $M_{\mathrm{src}}$ as:
$M=M_{\mathrm{src}}(1+z_{\mathrm{cos}})(1+z_{\mathrm{dop}})\mbox{\quad,\quad}$ (1)
where $z_{\mathrm{cos}}$ is the cosmological redshift, and $z_{\mathrm{dop}}$ is the Doppler shift caused by any constant line of sight (LOS) velocity (LOSV) of the CoM of the binary. A constant LOSV is fully degenerate with the source frame mass. However, a time-varying LOSV results in a time-varying Doppler shift and induces a time-varying detector frame mass, allowing time derivatives of the LOSV to be measured. Modulations to the GW waveform appear at different PN orders depending on the order of the time derivative of LOSV causing the modulation.
$M_{z}=M\left(1+\sum_{n=1}^{\infty}\Gamma_{n}(t_{\rm o}-t_{\rm c})^{n}\right)$ (2)
where $\Gamma_{n}\equiv\left.\frac{1}{c}\frac{1}{n!}\frac{d^{n}v_{l}}{dt_{\rm o}^{n}}\right|_{t_{\rm o}=t_{\rm c}}$, $t_{\rm o}$ is the observation time, $t_{\rm c}$ is the time at coalescence, and $c$ is the speed of light. Eq. (2) is valid under the assumption that $|\Gamma_{n}(t_{\rm o}-t_{\rm c})^{n}|$ are $\ll$ 1. As a consistency check of this condition, we ensure that $|\Gamma_{n}(t_{\rm o}-t_{\rm c})^{n}|\leq 0.05$, where special relativistic effects may be neglected.
The corresponding modulated inspiral GW waveform, in the stationary phase approximation, may be written as Buonanno et al. (2009):
@@ II.1 Phase Correction
where $\mathcal{A}\propto\sqrt{\frac{5}{24}}\frac{\mathcal{M}^{5/6}}{D_{\rm L}}$ is the amplitude, $\mathcal{M}$ is the chirp mass, $D_{\rm L}$ is the luminosity distance, $\Psi(f)$ is the unperturbed phase given by Eq. 3.18 of Buonanno et al. (2009), and $\Delta\Psi_{-4n}(f;\Gamma_{n})$ is a 4n-4n PN phase correction due to $n^{th}$ derivative of $v_{l}$ and at the leading order, it is given by:
@@
where $v=(\pi Mf)^{1/3}$, $\eta\equiv m_{1}m_{2}/(m_{1}+m_{2})^{2}$ is the symmetric mass ratio, and $f$ is the observed GW frequency. We consider the CBCs to be quasi-circular and non-spinning. Although CBCs at low frequencies are expected to have a residual eccentricity, the corresponding leading order LOSA phase corrections Samsing et al. (2024) are identical to the quasi-circular case.
Positive powers of a kinematic parameter can contribute to PN orders lower than the corresponding leading order. For example, the 8-8 PN phase correction $\Delta\Psi_{-8}(f)$ will have contributions from the LOS jerk (LOSJ) $\Gamma_{2}$ of the CoM, as well as the LOSA term proportional to $\Gamma_{1}^{2}$. We account for these corrections as well, down to 24-24 PN (see Appendix A.2), below which we do not expect PN terms to be well constrained even with detectors sensitive down to mili-hertz frequencies (see Figure 10 of the Appendix D).
For completeness, in Appendix B, we also compute the amplitude corrections. However, we do not use them while calculating the Fisher Matrix. We do not expect this choice to impact our results, and we show this in Appendix D by considering one of the specific cases of this analysis. In particular, we show that the leading order (relative) contributions to the amplitude due to the derivatives of LOSV are small i.e., $\ll 1$ (see the right panel of Figure 10 in Appendix D).
@@
II.2 Kinematics of the CoM of the CBC
To analyse the kinematics of the CoM of the CBC in these environments, we set up a Cartesian coordinate system in Figure 1 placed at the center of the gravitating source. Let $\iota_{\mathrm{out}}$ be the angle between our LOS and the angular momentum vector of the outer orbit, and let the projection vector of our LOS onto the outer orbital plane be along $\bm{e}_{y}$. Then:
@@ II.2 Kinematics of the CoM of the CBC
where $\cdot\equiv\frac{d}{dt}$ and $^{\prime}\equiv\frac{d}{dr}$. Using this relation iteratively, we calculate different kinematic parameters of the CoM of the binary and then project them onto our LOS ($\bm{n}$), which can be written as:
$\dot{\mathcal{V}}_{l}\equiv\dot{\bm{\mathcal{V}}}\cdot\bm{n}=[(\mathcal{V}^{\prime}_{r}\dot{r}-\mathcal{V}_{\vartheta}\dot{\vartheta})\sin\vartheta+(\mathcal{V}^{\prime}_{\vartheta}\dot{r}+\mathcal{V}_{r}\dot{\vartheta})\cos\vartheta]\sin\iota_{\mathrm{out}}.$ (8)
We provide the derivation as well as the mathematical expressions of the LOS components in the Appendix C.
@@
The $\Gamma_{n}$’s will, in general, be functions of environment and outer-orbit parameters, such as the steepness of the gravitational potential profile, and the location of the CBC with respect to the center of the potential, at coalescence. For instance, a circular orbit of radius $R$ in a Keplerian potential around an SMBH of mass $M_{\rm SMBH}$ would have $\Gamma_{n}\sim\omega^{n+1}R$, where $\omega^{2}\sim M_{\rm SMBH}/R^{3}$ (see Appendix C.1).
II.3 Fisher Matrix Analysis
@@
To assess our ability to constrain the environment and outer-orbit parameters in various observing scenarios, we adopt a Fisher Matrix approach Cutler and Flanagan (1994). We construct the Fisher matrix in $\{\ln D_{\rm L},\ln\mathcal{M},\ln\eta,\Gamma_{1},\Gamma_{2}\}$. We add or remove $\Gamma_{3}$, $\Gamma_{4}$, $\Gamma_{5}$, and $\Gamma_{6}$ to this set, as required, depending on the number of parameters that characterize the environment and outer-orbit. The PN corrections to the GW phase due to the kinematic parameters are rewritten in terms of the environment and outer-orbit parameters. The Fisher matrix is then inverted to yield a Covariance Matrix, which provides the Gaussian approximation to the GW likelihood. The environment parameters are sampled from this likelihood using emcee Foreman-Mackey et al. (2013), and the relative r.m.s errors on these parameters are estimated.
While calculating the covariance matrix, we re-write Eq. (2.7) of Cutler and Flanagan (1994) as:
$\Gamma_{ij}=4\mathrm{Re}\int_{f_{l}}^{f_{h}}\left(\frac{\partial\ln\tilde{h}(f)}{\partial\theta_{i}}\right)^{\ast}\frac{\partial\ln\tilde{h}(f)}{\partial\theta_{j}}|\tilde{h}(f)|^{2}\frac{df}{S_{n}(f)}$
where $f_{l}$ and $f_{h}$ are the frequencies when the signal enters the detector band and when it leaves the band, respectively. To avoid dealing with large numbers and numerical over/under-flows, let us define $\nu\equiv f/f_{0}$ and $S(\nu f_{0})\equiv S_{n}(\nu f_{0})/S_{0}$. Then using Eq. (3), we can write the above Eq. as $\Gamma_{ij}=4|\mathcal{A}|^{2}\frac{f_{0}^{-4/3}}{S_{0}}I_{ij}$ where:
$I_{ij}=\mathrm{Re}\int_{\nu_{l}}^{\nu_{h}}\left(\frac{\partial\ln\tilde{h}(\nu f_{0})}{\partial\theta_{i}}\right)^{\ast}\frac{\partial\ln\tilde{h}(\nu f_{0})}{\partial\theta_{j}}\frac{\nu^{-7/3}}{S(\nu f_{0})}d\nu$ (9)
$\nu_{l}=f_{l}/f_{0}$, and $\nu_{h}=f_{h}/f_{0}$. We can then write the covariance matrix as:
$C=\bm{\Gamma}^{-1}=\frac{S_{0}f_{0}^{4/3}}{4|\mathcal{A}|^{2}}\bm{I}^{-1}$ (10)
The quantities $f_{0}$ and $S_{0}$, in general, can be chosen to be any number that makes the inversion of the matrix $\bm{I}$ precise. However, here we choose them to be (roughly) the frequency at which the detector is most sensitive, and the corresponding value of the PSD, respectively. We further assume $4$ years of observation time for DECIGO and LISA, and the full sensitivity band of A+ and ET. Following Berti et al. (2005), we then set a sensitivity band of $10^{-2},10]\,{\rm Hz}$ ($10^{-4},1]\,{\rm Hz}$) for DECIGO Yagi and Seto (2011, 2017) (LISA Robson et al. (2019)), while for A+ and ET, we use the PSDs provided in LIGO Scientific Collaboration (2020) and Einstein Telescope Collaboration (2021); Hild et al. (2011), respectively.
The kinematic parameters in (A) are in the observer’s frame while the relations derived in the subsections (C.1 - C.4) are in the source (unperturbed) frame. For a cosmological redshift $z_{\rm cos}$, the source frame time and observer frame time are related as $t_{u}=t_{\rm o}/(1+z_{\rm cos})$. This leads to the rescaling of the kinematic parameters $\Gamma_{n}$ by a factor [^1] It can be seen straightforwardly that $\Gamma_{n,s}t_{s}^{n}=\Gamma_{n,s}\frac{t_{\rm o}^{n}}{(1+z_{\rm cos})^{n}}=\Gamma_{n,o}t_{\rm o}^{n}$ where $\Gamma_{n,o}\equiv\Gamma_{n,u}/(1+z_{\rm cos})^{n}$ $1/(1+z_{\rm cos})^{n}$. Therefore, strictly, estimating errors on the source-frame kinematic parameters will need to incorporate errors on the estimation of $z_{\rm cos}$, which can be acquired from the Fisher matrix analysis posterior on $D_{\rm L}$ and assuming standard cosmology. However, since we consider CBCs at low redshifts ($\sim 0.022$ and $\sim 0.198$ corresponding to $100\,{\rm Mpc}$ and $1\,{\rm Gpc}$, respectively), we neglect errors on $z_{\rm cos}$, i.e., assume it to be perfectly measured, and use the rescaled $\Gamma_{n}$’s in calculations as well as during the sampling with true value of $z_{\rm cos}$.
II.4 Signal-to-Noise Ratio (SNR)
The optimal SNR is given by:
$\varrho=\sqrt{4|\mathcal{A}|^{2}\mathcal{\int}_{f_{l}}^{f_{h}}\frac{f^{-7/3}}{S_{n}(f)}df}$ (11)
where $f_{l}$ and $f_{h}$ are the detector bandwidth’s lower and upper frequencies. $S_{n}(f)$ is the noise power spectral density (PSD). We assume face-on inner orbits which maximizes the SNR. When the binary is inclined, the SNRs will be reduced by a factor $\frac{\mathcal{Q}(\iota)}{\mathcal{Q}(\iota=0)}$ where:
$\mathcal{Q}(\iota)=\sqrt{\left(\frac{1+\cos^{2}\iota}{2}\right)^{2}+\cos^{2}\iota}$ (12)
and $\iota$ is the inclination of the (inner) binary relative to the LOS Robson et al. (2019).
Figure 2: Circular outer-orbit around an SMBH: The precision of the extraction of parameters over a grid of SMBH masses $M_{\rm SMBH}$ and radii of outer orbit $R$ for a binary system of component masses (from left to right) $5-1.4\,{\rm M}_{\odot}$ at $\mathrm{100Mpc}$ in A+, $30-30\,{\rm M}_{\odot}$ at $100\mathrm{Mpc}$ in ET, and $100-100\,{\rm M}_{\odot}$ at $1\mathrm{Gpc}$ in DECIGO and LISA, with $\cos\vartheta=0.9$. Top Panels: Relative error in the extraction of mass of SMBHs. The ‘solid’ lines are the lines of constant $\frac{\Delta M_{\rm SMBH}}{M_{\rm SMBH}}$. Bottom Panel: Relative error in the extraction of radii of the outer orbit. The ‘dashed’ lines are the lines of constant $\frac{\Delta R}{R}$. The pink regions correspond to the parameter space where either the relative errors are greater than 1 or we do not sample them due to the non-measurability of the kinematic parameters. The blue lines demarcate the regions (lower left) where we do not perform the Fisher matrix analysis, for, in these regions, the assumption $|\Gamma_{n}(t_{\rm o}-t_{\rm c})^{n}|\ll 1$ is no longer valid. The arrows point towards the region of increasing precision in the inference of parameters.
III Results